OpenStudy (anonymous):

55% of what number is 70?

OpenStudy (anonymous):

translating that into mathematical symbols you have: \(\large .55 \cdot x = 70 \) now just solve for x....

Parth (parthkohli):

\[55\% \times x = 70\]\[\implies 0.55 \times x = 70\]

OpenStudy (anonymous):

u all say the same thing :P

OpenStudy (unklerhaukus):

\[55\% \times n=70\] \[\frac{55}{100}n=70\] \[n=70\times\frac{100}{55}\]

OpenStudy (unklerhaukus):

\[n=70\times\frac{100}{55}=2\times5\times7\times \frac{2^2\times5^2}{5\times11}=\quad \dots\]